# Foundation Models in Medical Imaging - A Review and Outlook

### v3

@@ 4.2.1 Generalist Vision-Language Foundation Models
PubMedCLIP VLM 2D Radiologyc Resnet50 + ViT-B/32 ITC 80K
UniMiSS VM CT(3D), X-Ray MiT MIM 15K
BiomedCLIP VLM Multimodal b ViT-B/16 + PubMedBERT ITC 15M
PMC-CLIP VLM 2D Radiology c ResNet50 + PubmedBERT ITC + MLM 1.65M
MEDVInT VLM Multimodal b ResNet50 + Transformer MLM 227K
LLaVA-Med VLM Multimodalb LLaVA ITC 60K
LVM-Med VM CT(2D), MRI, X-ray, US ResNet-50 + ViT Graph Matching 1.3M
Med-Flamingo VLM
Multimodalb
ViT/L-14 + LLaMA-7B ITC 1.6M
RadFM VLM X-Ray, CT & MRI (3D), US, PET, MG 3D ViT + MedLLaMA-13B Generative ITC 16Md
Qilin-Med-VL VLM
Multimodalb
ViT/L-14 + LLaMA-13B ITC 580K
SAT VLM
Multimodalb
3D U-Net + BioBERT ITC 302K
VISION-MAE VM CT, MRI, PET, X-Ray, US Swin-T MAE 2.5M
RadCLIP VLM CT (2D + 3D), X-Ray, MRI ViT-L/14 Contrastive 1.2M
  • a
    X-Ray, US, MRI, PET, fluoroscopy, mammography, angiography
  • b
    Multimodal denotes broad range of radiological imaging types, as well as pathological image types.
  • c
    Wide range of 2D radiological imaging types, such as CT, MRI, PET, X-Ray, US, fMRI, ENG, Radioisotope, Endoscope, Mitotic, DOT and Fluorescence.
  • d
    2D image-text pairs, including 15.5M 2D images and 500k 3D images.
4.2.2 Chest X-Ray Vision-Language Foundation Models
One subdomain within radiology that has seen relatively more FM related developments than other modalities is Chest X-ray (CXR). CXR lends itself well to foundation model research due to its widespread use in clinical settings, the availability of large and well-established datasets such as [124, 125], and its relatively reduced computational complexity compared to more intricate 3D modalities like CT and MRI. However, unlike the earlier mentioned generalist models that can scale by incorporating multiple imaging modalities, gathering large, paired CXR-specific datasets remains challenging. As a result, some models like MedCLIP [126] have addressed this by leveraging unpaired data with semantic matching, enabling the use of more diverse datasets. Other models focused on enhancing the quality of existing paired data to maximize the learning potential and improve feature extraction. For example, CheXzero [127] employed a contrastive learning approach using only the concise "impressions" section of radiology reports, leading to better feature extraction and zero-shot disease classification and CXR-CLIP [128] built on this by creating high-quality image-text pairs with radiologist-designed prompts, further improving feature retrieval and interpretability. KAD [129] incorporated a medical knowledge graph into its training process, allowing for deeper reasoning and improved visual representation learning, although its reliance on single-source data limited its adaptability. To overcome this, UniChest [130] introduced a conquer-and-divide framework that systematically captured shared patterns across diverse datasets while isolating source-specific variations, resulting in better generalization across CXR benchmarks. Similarly, MedKLIP [131] focused on fine-grained alignment between images and text to enhance interpretability and disease classification, while MaCo [132] introduced a masked contrastive approach, refining image-text alignment and further improving zero-shot classification performance.
In addition to knowledge-enhancing models, others like ELIXR [133] and MAIRA-1 [134] focused on efficient adaptation by using specialized adapters to align visual encoders with large language models, reducing the need for full retraining. A more recent adaptation of MAIRA-1, MAIRA-2, combines two pre-trained image- and language encoders and adapts them for report generation from chest X-ray images [135]. More recently, Google’s CheXAgent [136] marked a shift toward scaling with large datasets. By leveraging the CheXinstruct dataset, which includes 6 million instruction-image-answer triplets, CheXAgent focused on large-scale instruction-tuned learning and demonstrated superior performance across multiple clinically relevant tasks, highlighting the potential of large-scale, instruction-based approaches.
Table 5: Foundation Models specialized for Chest X-Ray
Model Type Backbone SSL Objective Data Size Availability
cmidrule6-7 Weights Data
BioViL-T VLFM Custom CNN–Transformer + BERT Temporal Multi-Modal 174.1K
ELIXR VLFM SupCon + T5 ITC 220K
MaCo VLFM ViT-B/16 + BERT MLM 377K
CXR-CLIP VLFM ResNet50 + BioClinicalBERT ITC 15M
UniChest VLFM ResNet50 + BioClinicalBERT ITC 686K
RAD-DINO VFM ViT-B/14 DinoV2 838K ✓/✗
CheXagent VLFM Ensemble ITC + IC 6.1Ma b b
Ray-DINO VFM ViT-L DinoV2 863K
  • a
    Does not provide pre-training dataset size, but finetunes on 6.1M samples.
  • b
    Planned to be released.
@@ 4.2.3 Extending to 3D Modalities
Unlike X-ray imaging, which is inherently 2D, CT and MRI are 3D modalities that present challenges for foundation models. Early efforts to apply vision-language models to these modalities often relied on 2D slices from 3D scans due to limitations in available datasets, such as ROCO, which excluded 3D volumes. CLIP-LUNG [137] addressed CT imaging but worked with 2D slices, missing crucial spatial context. Subsequent models like M3FM [138] and MedBLIP [139] began to incorporate 3D data by dividing scans into patches or sub-volumes. However, these methods still did not fully leverage the spatial richness of 3D volumes, highlighting a gap in the development of fully integrated 3D foundation models for radiology.
@@
Although research into 3D foundation models for radiology is not at the level of 2D yet, significant progress was made with RadFM [140], which processed full 3D volumes using a 3D vision encoder and introduced the MedMD dataset, containing 16M 2D and 3D image-text pairs. Despite this advancement, RadFM faced limitations in vision-specific tasks like segmentation, partly due to the predominance of 2D images in its dataset. Recent models like M3D-LaMed [141] and SAT [142] have made strides in 3D segmentation, with M3D-LaMed introducing a novel 3D dataset and focusing on CT imaging, while SAT advanced segmentation capabilities across both CT and MRI. Further advancements include Rad-CLIP [143] and CT-CLIP [144], which extended the CLIP architecture to 3D imaging. Rad-CLIP relied on pseudo-3D structures from 2D slices, whereas CT-CLIP used a 3D vision transformer and a dataset of 50K 3D chest CT volumes. Merlin [145] represents a notable step forward in handling 3D CT data, integrating structured EHR data and unstructured text. However, Merlin’s focus on abdominal CT underscores a limitation in its applicability as a general foundation model. Overall, while these models demonstrate significant progress, they also reveal that comprehensive 3D foundation models for radiology are still in development.
Table 6: Foundation Models specialized for CT and/or MRI. Dataset size for pre-training is given in number of 3D volumes, unless stated otherwise.
Model Type Imaging Domain Backbone SSL Objective Data Size Availability
Weights Data
DeSD VM CT 3D ResNet50 DeSD 11K
SMIT VM CT, MRI SWIN-small MIM + Self-Distillation 3643
Medical Transformer VM MRI ResNet-18 + Transformer MAE 1783
M3FM VLM CT 3D CT-ViT + Transformer ITC 163K
CLIP-LUNG VLM CT ViT-B/16 + ResNet18 ITC 1010
Niu et al. VM CT 3D ViT Region-Contrastive 684
MedBLIP VLM MRI MedQFormer + BioMedLMa ITC 30K
MeTSK VM fMRI STGCN GraphCL 1415
Pai et al. VM CT 3D ResNet Modified SimCLR 11.5K
M3D-LaMed VLM CT 3D ViT + LLaMA-2 ITC 120K
Merlin VLM CT I3D ResNet152 + Clinical-Longformer Contrastive 25K a a
CT-CLIP VLM CT 3D CT-ViT + CXR-Bert ITC 26K
  • a
    Planned to be released.
  • b
@@
    Authors experiment with multiple language encoders. Table shows their best performing model architecture.
@@ 4.2.4 Vision-Only Models
While vision-language models have been predominant in radiology, driven by tasks like report generation and question answering, there is growing interest in developing vision-only foundation models. VFMs are better suited for spatially focused tasks such as detection and segmentation, which require precise understanding of the (3D) structure in medical imaging. For instance, Medical Transformer [146] and research by [147] employed multi-view and sequence-based approaches to handle 3D data, though these models only partially leveraged the full spatial potential of 3D volumes.
As the field progressed, various self-supervised learning techniques emerged. Initial approaches, such as those using contrastive learning, were inspired by vision-language models. Later, knowledge distillation techniques gained prominence, with models like DeSD [148] and SMIT [149] improving 3D segmentation through self-distillation and masked image modeling. The framework introduced by UniMiSS [150] expanded on these concepts, accommodating both 2D and 3D data. Additionally, more recent models explored graph-based SSL methods, such as LVM-Med [151] and MeTSK [152], which use graph structures to enhance feature representation across different imaging modalities.
Some works also have started using the DINOv2 framework for medical image analysis tasks. For instance, RAD-DINO [153], adapted the DINOv2 technique for medical imaging, showing strong performance in image registration and disease classification. Similarly, [154] apply DINOv2 to x-ray data. Their model, RayDINO, is among the first VFMs developed specifically for X-ray imaging. Trained on 873,000 X-rays, RayDINO adopts a comprehensive approach by not relying on assumptions about predefined classes. Instead, it focuses on extracting general features from imaging data, enabling it for example to detect rare diseases that may not be well-represented in training data. Compared to CheXzero, KAD and UniChest, which rely on textual supervision, RayDINO surpasses them in tasks such as multi-label disease detection, segmentation, and even report generation, while also generalizing well to unseen populations, achieving these results solely through large-scale self-supervised learning on imaging data.
Further advancements include models like VISION-MAE [155], which applied masked autoencoders to 3D data but was limited by its slice-level processing. [156] introduced a cancer-focused model trained on 3D CT data, excelling in biomarker discovery despite its narrower focus. The Segment Anything Model (SAM) has been tested across various studies, including [157], [158], [159], and [160], demonstrating its versatility in medical image segmentation while also facing challenges in cross-domain adaptation and efficient deployment.
@@ 5 Ophthalmology
Table 7: Summary of FMs for ophthalmology. Type refers to whether the FM is a vision model (VM) or a vision-language model (VLM). Modality refers to the type of ophthalmological data the model pre-trained on.
Model Type Imaging Domain Backbone Self-Supervised Learning Size Availability
Data Model Code Data
@@
  • a
    All datasets are publicly available except for the MIDAS dataset, which is subject to controlled access via an application process.
  • b
    All datasets are publicly available except for one private MRI dataset.
  • c
    CFP, FFA, OCTA, OCT, Slit-Lamp, B-Scan Ultrasound.
  • d
    150K real and 1M synthetic generated images.
  • e
    CFP, FFA, ICGA, FAF, RetCam, Ocular Ultrasound, OCT, Slit-Lamp, External Eye Photo, Specular Microscope, Corneal Topography.
  • f
    Planned to be released.
  • g
    No architectural details available.
5.1 Foundation Models for Ophthalmology
5.1.1 Foundation Models for Fundus Photography and OCT
The first foundation model for ophthalmology was RETFound by [161]. RETFound has a ViT model as its backbone, which is pre-trained on a large-scale dataset of over a million unlabelled retinal images using MAE during self-supervised learning. Two models are pretrained, either on OCT data only or CFP data only. After pre-training, these RETFound models are fine-tuned for specific disease detection tasks using explicit labels. The models can identify and quantify diseases such as diabetic retinopathy and AMD. Furthermore, with minimal fine-tuning, they could also detect other conditions, such as heart failure, stroke, and Parkinson’s disease.
Expanding on RETFound, [162] propose RETFound-Green, a variant of the original RETFound CFP model that focuses on efficiency. The model utilizing a novel token reconstruction pre-training objective, that helps it learn to identify features that are important for disease detection. This method enabled training with significantly less data and compute resources (as seen in Table 7) while maintaining comparable performance. Furthermore, DERETFound by [163] enhances performance by integrating stable diffusion techniques to generate synthetic data and incorporating medical expertise through text-tagging of images. This approach achieved comparable or superior results to RETFound using a fraction of the original dataset.
Another FM for ophthalmology that integrates textual expert-knowledge is FLAIR [164], which uses CLIP to align image features from fundus photography with textual descriptions. This integration of expert knowledge has shown to enhance performance in diabetic retinopathy grading, glaucoma detection, and multiclass disease diagnosis [165]. Another model built for CFP data is DRStageNet [166], which unlike the more generalist models described so far, is tailored specifically for diabetic retinopathy staging. DRStageNet employs a ViT backbone and DinoV2 for training. Its specialized architecture has demonstrated high accuracy and consistency in clinical settings.
@@ 5.1.2 Multi-Imaging Ophthalmology Foundation Models
While previous models were exclusively trained on fundus photography or OCT, VisionFM [167] integrates a variety of ophthalmic imaging modalities, including angiography, slit-lamp, and ultrasound. These imaging techniques can be used for visualizing other parts of the eye, not just retina, but also iris or cornea amongst others. VisionFM can diagnose diseases across these modalities using a single decoder, supporting a broad range of tasks like lesion segmentation and landmark detection, while generalizing efficiently to new imaging types. However, the paper does not yet provide specifics about the model architecture or training method, leaving the full scope of the results unclear at present.
More recently, EyeFound [168] was introduced as a general ophthalmic imaging FM, featuring a large ViT backbone and MAE pre-training, initialized with RETFound’s weights. The model is trained on a dataset of 2.8 million images from diverse ophthalmic imaging modalities, as specified in Table 7. EyeFound achieves higher AUROC scores for glaucoma detection and DR evaluation, and it outperforms RETFound in multi-label disease classification and systemic disease prediction. Additionally, it is evaluated on a visual question-answering task and demonstrates faster convergence during fine-tuning. However, its performance is not universally superior; in some datasets, its diagnostic accuracy is comparable to RETFound, and in tasks like myocardial infarction prediction, it does not significantly surpass RETFound, indicating areas for improvement in handling diverse and imbalanced datasets.
@@ 5.1.3 Techniques for Scaling Data
Beyond these FMs, smaller-scale models have introduced novel methods to address the problem of data availability in retinal image analysis. [169] propose FundusNet, a Resnet50 model pre-trained through contrastive learning with Neural Style Transfer (NST) augmentations. NST enhances training data diversity by transferring image styles, which is then used in a contrastive learning framework to train the model to maximize the similarity of different views of the same image and minimize the similarity of different images. This approach allows FundusNet to achieve improved performance and generalization across different datasets.
Furthermore, [170] focus on annotation-efficient learning for OCT segmentation. Their method uses generative self-supervised learning to train a ViT model on OCT images, which is further fine-tuned with a CNN encoder head for dense pixel-wise predictions. To reduce the annotation workload, they develop a selective annotation algorithm based on the greedy approximation for the k-center problem, constructing a representative subset of the data for annotation and significantly reducing the annotation cost.

### v4

@@ 4.2.2 Chest X-Ray Vision-Language Foundation Models
BioViL-T VLM Custom CNN–Transformer + BERT Temporal Multi-Modal 174.1K
ELIXR VLM SupCon + T5 ITC 220K
MaCo VLM ViT-B/16 + BERT MLM 377K
CXR-CLIP VLM ResNet50 + BioClinicalBERT ITC 15M
UniChest VLM ResNet50 + BioClinicalBERT ITC 686K
RAD-DINO VM ViT-B/14 DinoV2 838K ✓/✗
CheXagent VLM Ensemble ITC + IC 6.1Ma b b
Ray-DINO VM ViT-L DinoV2 863K
(a) Does not provide pre-training dataset size, but finetunes on 6.1M samples. (b) Planned to be released.
@@ 4.2.3 Extending to 3D Modalities
Unlike X-ray imaging, which is inherently 2D, CT and MRI are volumetric modalities that pose additional challenges for foundation models. Most FM architectures, such as vision transformers and contrastive vision–language frameworks, are optimized for 2D images. Applying them directly to 3D data requires either slicing volumes into 2D images, thereby losing important spatial context, or developing dedicated 3D encoders, which face obstacles including higher computational cost, limited availability of large-scale 3D datasets, and the difficulty of learning rich volumetric representations. As a result, early radiology FMs often defaulted to processing 3D scans as collections of 2D slices.
Early examples include CLIP-LUNG [lei2023clip], which adapted the CLIP framework to CT but only at the slice level, thereby missing volumetric information. Subsequent works, such as M3FM [niu2023medical] and MedBLIP [chen2023medblip], began to incorporate 3D data by dividing scans into 3D patches or sub-volumes, though these approaches still offered only partial use of spatial context.
@@
More recent models have advanced toward fully volumetric approaches. RadFM [wu_towards_2023] was a notable milestone, introducing the MedMD dataset (16M 2D and 3D image–text pairs) and employing a 3D vision encoder. However, performance remained limited for vision-specific tasks like segmentation, partly due to the dominance of 2D data in the training set. Other efforts, such as M3D-LaMed [bai2024m3d] and SAT [zhao_one_2023], focused specifically on 3D segmentation, while Rad-CLIP [lu2024radclip] and CT-CLIP [hamamci_foundation_2024] extended CLIP-style training to volumetric data. Among them, Rad-CLIP relied on pseudo-3D structures from 2D slices, whereas CT-CLIP used a 3D vision transformer and a dataset of 50K 3D chest CT volumes. Merlin [blankemeier2024merlin] represents a notable step forward in handling 3D CT data, integrating structured EHR data and unstructured text. However, its focus on abdominal CT limits its applicability as a general-purpose foundation model.
Taken together, these works illustrate a gradual shift from 2D slice-based training toward models capable of handling entire 3D volumes. While progress has been made, especially in segmentation and volumetric representation learning, the field still lacks large-scale, general-purpose 3D foundation models comparable to their 2D counterparts.
Table 6: Foundation Models specialized for CT and/or MRI. Type refers to whether the FM is a vision model (VM) or a vision-language model (VLM). Backbone and SSL objective show how the model was pretrained. Imaging domain shows which medical image types were used to pre-train the model. Dataset size for pre-training is given in number of 3D volumes, unless stated otherwise. Availability shows whether the weights and pretraining data are available online.
Model Type Imaging Backbone SSL Objective Data Size Availability Weights Data DeSD VM CT 3D ResNet50 DeSD 11K SMIT VM CT, MRI SWIN-small MIM + Self-Distillation 3643 Medical Transformer VM MRI ResNet-18 + Transformer MAE 1783 M3FM VLM CT 3D CT-ViT + Transformer ITC 163K CLIP-LUNG VLM CT ViT-B/16 + ResNet18 ITC 1010 Niu et al. VM CT 3D ViT Region-Contrastive 684 MedBLIP VLM MRI MedQFormer + BioMedLMa ITC 30K MeTSK VM fMRI STGCN GraphCL 1415 Pai et al. VM CT 3D ResNet Modified SimCLR 11.5K M3D-LaMed VLM CT 3D ViT + LLaMA-2 ITC 120K Merlin VLM CT I3D ResNet152 + Clinical-Longformer Contrastive 25K a a CT-CLIP VLM CT 3D CT-ViT + CXR-Bert ITC 26K
@@
(a) Planned to be released. (b) Authors experiment with multiple language encoders. Table shows their best performing model architecture.
@@ 4.2.4 Vision-Only Models
While vision-language models have been predominant in radiology, driven by tasks like report generation and question answering, there is growing interest in developing vision-only foundation models. VFMs are better suited for spatially focused tasks such as detection and segmentation, which require precise understanding of the (3D) structure in medical imaging. For instance, Medical Transformer [jun2021medicaltransformeruniversalbrain] and research by [niu_unsupervised_2022] employed multi-view and sequence-based approaches to handle 3D data, though these models only partially leveraged the full spatial potential of 3D volumes.
As the field progressed, various self-supervised learning techniques emerged. Initial approaches, such as those using contrastive learning, were inspired by vision-language models. Later, knowledge distillation techniques gained prominence, with models like DeSD [ye_desd_2022] and SMIT [jiang_self-supervised_2022] improving 3D segmentation through self-distillation and masked image modeling. The framework introduced by UniMiSS [xie2022unimiss] expanded on these concepts, accommodating both 2D and 3D data. Additionally, more recent models explored graph-based SSL methods, such as LVM-Med [nguyen_lvm-med_2023] and MeTSK [cui_meta_2023], which use graph structures to enhance feature representation across different imaging modalities.
Some works also have started using the DINOv2 framework for medical image analysis tasks. For instance, RAD-DINO [perez-garcia_rad-dino_2024], adapted the DINOv2 technique for medical imaging, showing strong performance in image registration and disease classification. Similarly, [moutakanni2024advancing] apply DINOv2 to x-ray data. Their model, RayDINO, is among the first VFMs developed specifically for X-ray imaging. Trained on 873,000 X-rays, RayDINO adopts a comprehensive approach by not relying on assumptions about predefined classes. Instead, it focuses on extracting general features from imaging data, enabling it for example to detect rare diseases that may not be well-represented in training data. Compared to CheXzero, KAD and UniChest, which rely on textual supervision, RayDINO surpasses them in tasks such as multi-label disease detection, segmentation, and even report generation, while also generalizing well to unseen populations, achieving these results solely through large-scale self-supervised learning on imaging data.
Further advancements include models like VISION-MAE [liu2024visionmae], which applied masked autoencoders to 3D data but was limited by its slice-level processing. [pai_foundation_2024] introduced a cancer-focused model trained on 3D CT data, excelling in biomarker discovery despite its narrower focus. The Segment Anything Model (SAM) has been tested across various studies, including [yan2022sam], [mazurowski2023segment], [zhang2023customized], and [gao_desam_2023], demonstrating its versatility in medical image segmentation while also facing challenges in cross-domain adaptation and efficient deployment.
@@ 5 Ophthalmology
5.1 Foundation Models for Ophthalmology
5.1.1 Foundation Models for Fundus Photography and OCT
The first foundation model for ophthalmology was RETFound by [zhou_foundation_2023]. RETFound has a ViT model as its backbone, which is pre-trained on a large-scale dataset of over a million unlabelled retinal images using MAE during self-supervised learning. Two models are pretrained, either on OCT data only or CFP data only. After pre-training, these RETFound models are fine-tuned for specific disease detection tasks using explicit labels. The models can identify and quantify diseases such as diabetic retinopathy and AMD. Furthermore, with minimal fine-tuning, they could also detect other conditions, such as heart failure, stroke, and Parkinson’s disease.
Expanding on RETFound, [engelmann2024training] propose RETFound-Green, a variant of the original RETFound CFP model that focuses on efficiency. The model utilizing a novel token reconstruction pre-training objective, that helps it learn to identify features that are important for disease detection. This method enabled training with significantly less data and compute resources (as seen in Table 7) while maintaining comparable performance. Furthermore, DERETFound by [yan2024expertise] enhances performance by integrating stable diffusion techniques to generate synthetic data and incorporating medical expertise through text-tagging of images. This approach achieved comparable or superior results to RETFound using a fraction of the original dataset.
Table 7: FMs for ophthalmology. Type refers to whether the FM is a vision model (VM) or a vision-language model (VLM). Modality refers to the type of ophthalmological data the model pre-trained on. Backbone and SSL objective show how the model was pretrained. Data size shows the number of 2D images used for pre-training, while model size denotes the number of parameters used for the model. Availability shows whether the weights and pretraining data are available online.
Model Type Imaging Backbone SSL Objective Size Availability
Data Model Weights Data
@@ 5.1.1 Foundation Models for Fundus Photography and OCT
(a) All datasets are publicly available except for the MIDAS dataset, which is subject to controlled access via an application process. (b) All datasets are publicly available except for one private MRI dataset. (c) CFP, FFA, OCTA, OCT, Slit-Lamp, B-Scan Ultrasound. (d) 150K real and 1M synthetic generated images. (e) CFP, FFA, ICGA, FAF, RetCam, Ocular Ultrasound, OCT, Slit-Lamp, External Eye Photo, Specular Microscope, Corneal Topography. (f) Planned to be released. (g) No architectural details available.
Another FM for ophthalmology that integrates textual expert-knowledge is FLAIR [silvarodriguez2023foundation], which uses CLIP to align image features from fundus photography with textual descriptions. This integration of expert knowledge has shown to enhance performance in diabetic retinopathy grading, glaucoma detection, and multiclass disease diagnosis [silva2024importance]. Another model built for CFP data is DRStageNet [men2023drstagenet], which unlike the more generalist models described so far, is tailored specifically for diabetic retinopathy staging. DRStageNet employs a ViT backbone and DinoV2 for training. Its specialized architecture has demonstrated high accuracy and consistency in clinical settings.
@@ 5.1.2 Multi-Imaging Ophthalmology Foundation Models
While previous models were exclusively trained on fundus photography or OCT, VisionFM [qiu2023visionfm] integrates a variety of ophthalmic imaging modalities, including angiography, slit-lamp, and ultrasound. These imaging techniques can be used for visualizing other parts of the eye, not just retina, but also iris or cornea amongst others. VisionFM can diagnose diseases across these modalities using a single decoder, supporting a broad range of tasks like lesion segmentation and landmark detection, while generalizing efficiently to new imaging types. However, the paper does not yet provide specifics about the model architecture or training method, leaving the full scope of the results unclear at present.
More recently, EyeFound [shi2024eyefound] was introduced as a general ophthalmic imaging FM, featuring a large ViT backbone and MAE pre-training, initialized with RETFound’s weights. The model is trained on a dataset of 2.8 million images from diverse ophthalmic imaging modalities, as specified in Table 7. EyeFound achieves higher AUROC scores for glaucoma detection and DR evaluation, and it outperforms RETFound in multi-label disease classification and systemic disease prediction. Additionally, it is evaluated on a visual question-answering task and demonstrates faster convergence during fine-tuning. However, its performance is not universally superior; in some datasets, its diagnostic accuracy is comparable to RETFound, and in tasks like myocardial infarction prediction, it does not significantly surpass RETFound, indicating areas for improvement in handling diverse and imbalanced datasets.
@@ 5.1.3 Techniques for Scaling Data
Beyond these FMs, smaller-scale models have introduced novel methods to address the problem of data availability in retinal image analysis. [alam_contrastive_2023] propose FundusNet, a Resnet50 model pre-trained through contrastive learning with Neural Style Transfer (NST) augmentations. NST enhances training data diversity by transferring image styles, which is then used in a contrastive learning framework to train the model to maximize the similarity of different views of the same image and minimize the similarity of different images. This approach allows FundusNet to achieve improved performance and generalization across different datasets.
Furthermore, [zhang_annotation-efficient_2023] focus on annotation-efficient learning for OCT segmentation. Their method uses generative self-supervised learning to train a ViT model on OCT images, which is further fine-tuned with a CNN encoder head for dense pixel-wise predictions. To reduce the annotation workload, they develop a selective annotation algorithm based on the greedy approximation for the k-center problem, constructing a representative subset of the data for annotation and significantly reducing the annotation cost.